### v2

# ReMA: Learning to meta-think for LLMs with multi-agent reinforcement learning

Abstract
Recent research on Reasoning of Large Language Models (LLMs) has sought to further enhance their performance by integrating meta-thinking—enabling models to monitor, evaluate, and control their reasoning processes for more adaptive and effective problem-solving. However, current single-agent work lacks a specialized design for acquiring meta-thinking, resulting in low efficacy. To address this challenge, we introduce Reinforced Meta-thinking Agents (ReMA), a novel framework that leverages Multi-Agent Reinforcement Learning (MARL) to elicit meta-thinking behaviors, encouraging LLMs to think about thinking. ReMA decouples the reasoning process into two hierarchical agents: a high-level meta-thinking agent responsible for generating strategic oversight and plans, and a low-level reasoning agent for detailed executions. Through iterative reinforcement learning with aligned objectives, these agents explore and learn collaboration, leading to improved generalization and robustness. Experimental results demonstrate that ReMA outperforms single-agent RL baselines on complex reasoning tasks, including competitive-level mathematical benchmarks and LLM-as-a-Judge benchmarks. Comprehensive ablation studies further illustrate the evolving dynamics of each distinct agent, providing valuable insights into how the meta-thinking reasoning process enhances the reasoning capabilities of LLMs.
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in knowledge understanding and complex reasoning tasks (Chowdhery et al., 2023; Achiam et al., 2023; Anil et al., 2023; Dubey et al., 2024). The paradigm in developing LLM-based reasoning models is shifting from scaling training-time computation towards scaling test-time computation (Snell et al., 2024). Recent advancements, such as OpenAI-o1 (OpenAI, 2024), Deepseek R1 (DeepSeek-AI et al., 2025), and Gemini 2.0 Flash Thinking (DeepMind, 2025), have demonstrated that allowing LLMs to think before generating answers can significantly enhance performance and lead to the emergence of human-like reasoning patterns. These patterns like “Wait, hold on. Let me double-check” or “Let’s break this down.”, indicate that LLMs can develop a form of meta-thinking abilities that can generalize well to out-of-distribution (OOD) tasks (Xiang et al., 2025). Meta-thinking, also known as metacognitive skills (Flavell, 1979), is an ability traditionally considered uniquely human (Didolkar et al., 2024).
To cultivate meta-thinking patterns from LLMs themselves, recent construction-based supervised approaches leverage supervised finetuning on structured reasoning trajectories. Specifically, these methods sampling reasoning trajectories from predefined meta-thinking templates and then use supervised finetuning (SFT) or direct preference optimization (DPO) (Rafailov et al., 2023) to teach LLMs imitate these patterns (Qi et al., 2024; Yue et al., ; Xi et al., 2024; Yang et al., 2025; Muennighoff et al., 2025; Ye et al., 2025c). However, such methods lack sufficient flexibility for LLMs to explore suitable meta-thinking patterns. Thus, they often fail to generalize to out-of-distribution (OOD) problems, leading to unstable performance on unseen data (Kirk et al., ; Chu et al., 2025). Besides construction-based methods, R1-like single-agent reinforcement learning (SARL) has also been adopted for meta-thinking in reasoning (DeepSeek-AI et al., 2025; Xie et al., 2025). However, these SARL attempts typically rely on strong foundational models for easier exploration or extensive task-specific fine-tuning for stable training (Xu et al., 2025; Gandhi et al., 2025). Furthermore, SARL needs to learn meta-thinking and reasoning within a single forward pass, seeking to capture complex reasoning structures purely in an autoregressive manner (Xie et al., 2025). This can potentially lead to issues such as inefficient exploration as well as reduced readability and early convergence to local optima (DeepSeek-AI et al., 2025; Xiang et al., 2025).
Figure 1: Left: A construction-based method that fine-tunes LLMs using rejection sampling, searching among combinations of pre-defined templates. Middle: RL-from-base method learns to mix meta-thinking and detailed solution steps during training. Right: Our method ReMA separates the meta-thinking and reasoning steps in a multi-agent system, allowing the agents to explore efficiently and learn to collaborate.
To address these limitations, we introduce Reinforced Meta-thinking Agents (ReMA), a novel framework that leverages multi-agent reinforcement learning (MARL) to encourage LLMs to think about thinking. Our approach employs a multi-agent system (MAS) composed of a high-level meta-thinking agent, responsible for strategic oversight and instruction generation, and a low-level reasoning agent tasked with detailed executing processes based on provided guidance. We compare the inference process among the construction-based method, R1-like method, and ReMA in Fig 1. Since MAS distributes the exploration space of SARL into multiple agents, it enables each agent to explore more structurally and efficiently during training. Then we apply reinforcement learning iteratively on each agent with aligned reward functions. In this way, ReMA effectively balances the trade-off between generalization capability and exploration efficiency. As a result, they can learn to play the best of their role (either to meta-think or to follow instructions), at the present of the other agent. Unlike SFT primarily leads to memorization (Chu et al., 2025), these two agents dynamically adjust their parameters and finally induce effective and robust collaboration. Consequently, not only does ReMA alleviate the OOD issues, but it also stimulates more organized learning of meta-thinking due to the hierarchy design.
To our knowledge, we are the first to formally define the concept of a multi-agent meta-thinking reasoning process (MAMRP) and optimize it through (multi-agent) reinforcement learning. Furthermore, we conduct extensive experiments on both math reasoning and LLM-as-a-Judge tasks. Across three backbone pretrained models on these benchmarks, ReMA achieves the highest average performance under the same training dataset. Additionally, we conduct comprehensive ablations to illustrate the evolving dynamics of each distinct agent. For instance, we observed unexpected interaction patterns characterized by a role reversal under different reward settings. These findings can provide valuable insights into how the meta-thinking reasoning process enhances the reasoning capabilities of LLMs.
2 preliminaries
In this section, we outline the formulation of the vanilla reasoning process (Sec. 2.1) and the representative training methods (Sec. 2.2) along with the notation used throughout the paper.
2.1 Vanilla Reasoning Process (VRP)
The autoregressive decoding process of a language model can be formalized as a series of sequential token predictions, where the overall probability is equal to the product of the probabilities at each individual step. Given a language model $\pi_{\theta}$ parameterized by $\theta$, and a task prompt $\mathbf{x}=(x_{0},x_{1},\ldots,x_{N})$ of $N$ tokens, the vanilla reasoning process (VRP) directly generates a response $\mathbf{y}=(y_{0},y_{1},\ldots,y_{L})$ of $L$ tokens. The probability of generating $\mathbf{y}$ is
$\pi_{\theta}(\mathbf{y}|\mathbf{x})=\prod\limits_{i=0}^{L}\pi_{\theta}(y_{i}|x_{0},x_{1},\ldots x_{N},y_{0},\ldots,y_{i-1}).$
In the specific case where the response contains structured intermediate reasoning steps, the process is then known as chain-of-thought (CoT) (Wei et al., 2022), and the corresponding directed graph is as follows:
[…]
where $\mathbf{a}$ is the extracted final answer, which is included in the answer $\mathbf{y}$.
2.2 Training VRP: Supervised Learning and Reinforcement Learning
Recently, there has been growing interest in uncovering this kind of reasoning process through RL on strong base models such as DeepSeek-V3-Base (Liu et al., 2024) and Qwen2.5-Math-7B (Yang et al., 2024). The decoding strategy of LLM follows the vanilla reasoning process and can further be formulated as a deterministic token-level Markov Decision process (Wang et al., 2024a). The optimization loss function can be denoted as follows:
$\mathcal{L}(\theta)=\mathbb{E}_{(\mathbf{x},\mathbf{y}^{*})\sim\mathcal{D},\mathbf{y}\sim\pi_{\theta}}\left[-\mathcal{R}(\mathbf{y},\mathbf{y}^{*})\right].$ (2)
where $\mathcal{R}(\cdot,\cdot)$ represents a reward function comparing generated answer $\mathbf{y}$ with the golden answer $\mathbf{y^{*}}$ for any question $\mathbf{x}$ sampled from dataset $\mathcal{D}$. To compute the gradient of the loss function with respect to the policy $\pi_{\theta}$, the REINFORCE algorithm (Williams, 1992) applies the log-derivative trick. Specifically, it computes the gradient as follows:
$\nabla_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{(\mathbf{x},\mathbf{y}^{*})\sim\mathcal{D},\mathbf{y}\sim\pi_{\theta}}\left[-\nabla_{\theta}\log\pi_{\theta}(\mathbf{y}\mid\mathbf{x})\mathcal{R}(\mathbf{y},\mathbf{y}^{*})\right].$
Since strict online sampling and gradient calculation can be computationally expensive, proximal policy optimization (PPO) (Schulman et al., 2017) is widely adopted for its high training stability and sampling efficiency. The PPO loss function is:
$\mathcal{L}(\theta)=\mathbb{E}_{(\mathbf{x},\mathbf{y}^{*})\sim\mathcal{D},\mathbf{y}\sim\pi_{\text{old}}}\left[\min\left(\rho(\mathbf{y})A^{\pi_{\text{old}}},\text{clip}(\rho(\mathbf{y}),1-\epsilon,1+\epsilon)A^{\pi_{\text{old}}}\right)\right],$ (3)
where $\rho(\mathbf{y})=\frac{\pi_{\theta}(\mathbf{y|x})}{\pi_{\text{old}}(\mathbf{y|x})}$ denotes the ratio between current and the old policy, $A^{\pi_{\text{old}}}$ is the advantage of policy $\pi_{\text{old}}$, i.e., the extra reward over the average reward. REINFORCE++ (Hu, 2025a) combines the simplicity of the REINFORCE algorithm with the enhanced training stability of PPO. While REINFORCE++ builds upon the REINFORCE framework, it further improves performance by incorporating techniques from PPO such as gradient clipping, mini-batch updates, and advantage estimation.
However, RL on base LLMs that haven’t been well-aligned may suffer from issues like poor readability and language mixing, preventing researchers from verifying, understanding, and further developing their LLMs. And huge searching space makes efficient learning of meta-thinking daunting.
3 Method
In this section, we present Reinforced Meta-Thinking Agent (ReMA), a reinforcement learning method integrating meta-thinking into the reasoning process of LLM under multi-agent settings (Sec 3.1), then describe the bilateral learning process enabled by MARL (Sec 3.2), and finally discuss the significant impact that reward function design has on learning dynamics (Sec 3.3).
3.1 Deploying Meta-Thinking Reasoning Process (MRP) for LLMs
Beyond VRP mentioned in Section 2.1, recent studies (Muennighoff et al., 2025; Ye et al., 2025c) have shown that integrating meta-thinking behaviors in reasoning process can largely improve the accuracy of the final answers. Meta-thinking is broadly defined as the ability to monitor, control, and evaluate one’s cognitive processes (Flavell, 1979). Effective meta-thinking fosters self-awareness, flexible thinking, and adaptive problem-solving, enabling individuals to recognize errors, seek alternative solutions, and optimize performance. We define the Meta-thinking Reasoning Process (MRP) for both single- and multi-agents. In a single-agent setting, such a reasoning process can be formulated by
$\pi_{\theta}(\mathbf{y|x})=\pi_{\theta}(\mathbf{y|x,m})\cdot\pi_{\theta}(\mathbf{m|x}),$ (4)
where $\mathbf{m}=(m_{0},m_{1},\ldots,m_{H})$ is the output tokens of meta-thinking during solving a question, which calls LLM’s once and generates meta-thinking and following reasoning autoregressively. Such a process can be further extended to a multi-turn setting as well.
Based on this formulation, we define the meta-thinking reasoning process (MRP) below:
[…]
Previous works are mainly construction-based methods, i.e. sampling from predefined templates of MRP and then applying supervised fine-tuning (Yue et al., ; Xi et al., 2024; Zhao et al., 2024). These works typically fail to fully explore the exponentially large language space and are confined to predefined patterns, even when using efficient search algorithms (Yao et al., 2023; Feng et al., 2023; Snell et al., 2024). To solve these problems, we separate the process and assign different LLMs. These LLMs simulate the meta-thinking reasoning process through conversation-like interactions between two agents: one is a high-level meta-thinking agent, and the other is a low-level reasoning agent.
During a conversation, the high-level and low-level agents (namely, $\pi_{\theta_{h}}$ and $\pi_{\theta_{l}}$) act in an interleaving manner. The high-level agent serves as the meta-thinking process, summarizing and generating meta-thoughts based on the question and the history of interactions, while the low-level agent focuses on detailed problem-solving and execution, following instructions provided by its counterpart. The single-turn MRP problem within the single-agent scenario, as formally defined by Eq.4, can be extended to a multi-agent setting. We formulate the multi-agent meta-thinking reasoning process (MAMRP) as follows:
$\mathbf{y}\sim\pi_{\theta_{l}}(\mathbf{y|x,m})\cdot\pi_{\theta_{h}}(\mathbf{m|x}).$ (6)
Similarly, we present the process with a directed graph with separated agents:
[…]
To illustrate this procedure, we consider a representative example: The high-level agent first conducts a meta-thinking analysis of the task’s objective, appropriate solution strategies, and problem decomposition schema. This analysis is then passed to the low-level agent as a guidance for step-wise solutions. MAMRP can also be extended to a multi-turn setting: then in subsequent turns, the high-level agent can also refine or critique the low-level agent’s outputs. The process repeats until any predefined stopping criteria are met, such as reaching the maximum number of iterations or achieving a satisfactory solution.
Moreover, as a complex reasoning system, MAMRP provides various optimization opportunities in scaling inference-time computation. We leave further discussion of these aspects in Appendix A.
3.2 Training MAMRP: A Multi-agent RL Method
Figure 2: Comparison of Training Pipelines. Left: RL training of VRP and MRP. Right: MARL training in ReMA: the high-level agent is frozen while the low-level agent is trained using generated meta-thinking, execution results, and rewards. The low-level agent is frozen while the high-level agent is trained. This cycle repeats iteratively.
Rather than supervised finetuning on predefined execution pipelines, LLM(s) must possess the flexibility to explore freely and update rapidly. Therefore it is necessary to adopt RL training. To optimize the multi-agent system proposed in Sec 3.1, we frame the problem and objective function within the RL structure. First, we define a joint hierarchical policy with sequential decisions as follows:
$\pi_{(\theta_{h},\theta_{l})}(\mathbf{y}\mid\mathbf{x})=\pi_{\theta_{l}}(\mathbf{y}\mid\mathbf{x},\mathbf{m}\sim\pi_{\theta_{h}}(\cdot\mid\mathbf{x})),$ (8)
Let $R(\mathbf{y,y}^{*})$ denote the final reward, which serves as the objective function $J(\theta_{h},\theta_{l})$ for the joint policy:
$J(\theta_{h},\theta_{l})=\mathbb{E}_{\mathbf{x},\mathbf{y}^{*}}\mathbb{E}_{\mathbf{y}\sim\pi_{(\theta_{h},\theta_{l})}}R(\mathbf{y},\mathbf{y}*).$ (9)
During optimization procedure, the high-level policy $\pi_{\theta_{h}}$ and low-level policy $\pi_{\theta_{l}}$ aim to maximize their respective rewards independently. To enhance the overall system performance, such as the final success rate $R$, their individual reward functions are designed to enhance effective collaboration, ultimately leading to improved performance. The optimization goals for agents are:
$\theta_{h}^{*}$ $=\arg\max_{\theta_{h}}\mathbb{E}_{(\mathbf{x,y^{*}})\sim\mathcal{D},\mathbf{m}\sim\pi_{\theta_{h}},\mathbf{y}\sim\pi_{\theta_{l}^{*}}}\left[R_{h}(\mathbf{m,y,y}^{*})\right],$ (10)
$\theta_{l}^{*}(\theta_{h})$ $=\arg\max_{\theta_{l}}\mathbb{E}_{(\mathbf{x,y^{*}})\sim\mathcal{D},\mathbf{m}\sim\pi_{\theta_{h}},\mathbf{y}\sim\pi_{\theta_{l}}}\left[R_{l}(\mathbf{m,y,y}^{*})\right],$ (11)
where $R_{h}$ and $R_{l}$ are policies’ individual reward functions, including $R$ and regularization according to tasks and models, e.g., different format rewards. In this work, we conduct experiments to enhance MAMRP described in Equation 7 using the update method outlined in Equations 10 and 11. For the detailed design of our experiments, we’ll discuss different designs of reward functions in Sec 3.3. With the above objective functions, we apply multi-agent RL algorithms to enhance collaboration between the two models. Additionally, to stabilize the training process, we conduct prompt filtering as described in Cui et al. (2025): during each round of training, we filter out the questions that our models solve with a success rate less than $\varepsilon_{\min}$ or greater than $\varepsilon_{\max}$. Similar to Kirchner et al. (2024b), we train the models iteratively with RL algorithms as shown in the Algorithm 1. We illustrate the MAMRP inference procedure and the proposed training method in Fig. 2. We also provide an analysis of different loss functions in Appendix F.
3.3 Reward Function Designs
As described in Section 3.2, we update both high-level and low-level agents by assigning rewards based on the low-level policy output. Below, we outline several potential reward designs:
  • Correctness reward: For tasks with explicit ground truth, we assign rewards based on the correctness of the low-level agent’s output.
  • Format reward: For tasks that require a specific output format, we enforce adherence to the prescribed structure by providing a format reward.
  • To encourage the high-level agent to generate informative and unambiguous meta-thinking, and to stabilize the low-level outputs, we reward the high-level agent when the low-level agent produces consistent responses. Specifically, the consistency reward is defined as
    $R_{h}=\frac{\text{max occurrence of an answer}}{\text{total number of responses}}.$
The combination of different rewards enhances the overall performance of LLMs. Detailed reward designs are presented in Appendix E.
4 Experiments
To evaluate the effectiveness and efficiency of ReMA, we conduct experiments on challenging benchmarks for two types of tasks: mathematical reasoning and LLM-as-a-Judge, and we train and test different LLMs.
Table 1: Performance on in-distribution test sets and out-of-distribution test sets. We also report the improvement/degradation w.r.t. basic CoT performance(VRP). On average, ReMA outperforms all baselines. Particularly on out-of-distribution datasets, ReMA achieves the highest performance on most of the benchmarks.
(a) Performance on math benchmarks
[…]
Llama3 -8B -Instruct MATH500 30.80 33.40 (+2.60) 32.80 (+2.00) 33.80 (+3.00)
GSM8K 67.48 81.80 (+14.32) 79.68 (+12.20) 79.38 (+11.90)
AIME24 0.00 0.00 (+0.00) 3.33 (+3.33) 0.00 (+0.00)
AMC23 2.50 10.00 (+7.50) 12.50 (+10.00) 22.50 (+20.00)
Gaokao2023en 22.34 27.53 (+5.19) 23.38 (+1.04) 28.57 (+6.23)
Minerva Math 8.82 16.54 (+7.72) 18.01 (+9.19) 13.97 (+5.15)
Olympiad Bench 8.44 8.89 (+0.45) 9.33 (+0.89) 8.89 (+0.45)
Average 20.05 25.45 (+5.40) 25.58 (+5.53) 26.73 (+6.68)
Llama3.1 -8B -Instruct MATH500 50.80 50.20 (-0.60) 48.60 (-2.20) 53.20 (+2.40)
GSM8K 86.05 84.53 (-1.52) 85.37 (-0.68) 87.26 (+1.21)
AIME24 10.00 3.33 (-6.67) 6.67 (-3.33) 13.33 (+3.33)
AMC23 27.50 12.50 (-15.00) 30.00 (+2.50) 20.00 (-7.50)
Gaokao2023en 38.96 36.10 (-2.86) 37.14 (-1.82) 37.14 (-1.82)
Minerva Math 22.79 26.84 (+4.05) 25.37 (+2.58) 28.31 (+5.52)
Olympiad Bench 15.11 19.70 (+4.59) 15.70 (+0.59) 19.56 (+4.45)
Average 35.89 33.32 (-2.57) 35.55 (-0.34) 36.97 (+1.08)
Qwen2.5 -7B -Instruct MATH500 75.00 77.20 (+2.20) 76.40 (+1.40) 74.40 (-0.60)
GSM8K 92.04 91.36 (-0.68) 91.81 (-0.23) 90.60 (-1.44)
AIME24 6.67 6.67 (+0.00) 10.00 (+3.33) 20.00 (+13.33)
AMC23 47.50 50.00 (+2.50) 52.50 (+5.00) 57.50 (+10.00)
Gaokao2023en 56.62 54.81 (-1.81) 55.06 (-1.56) 57.92 (+1.30)
Minerva Math 35.66 34.93 (-0.73) 32.35 (-3.31) 34.93 (-0.73)
Olympiad Bench 38.22 38.37 (+0.15) 37.78 (-0.44) 36.30 (-1.92)
Average 50.24 50.48 (+0.24) 50.84 (+0.60) 53.09 (+2.85)
(b) Performance on LLM-as-a-Judge benchmarks
[…]
Llama3.1 -8B -Instruct RewardBench970 69.48 82.89 (+13.41) 81.13 (+11.65) 83.71 (+14.23)
JudgeBench 51.29 51.94 (+0.65) 52.90 (+1.61) 52.90 (+1.61)
Average 60.39 67.41 (+7.02) 67.02 (+6.63) 68.31 (+7.92)
Qwen2.5 -7B -Instruct RewardBench970 78.56 85.36 (+6.80) 86.49 (+7.93) 83.51 (+4.95)
JudgeBench 58.39 56.94 (-1.45) 58.39 (+0.00) 56.94 (-1.45)
Average 68.47 71.15 (+2.68) 72.44 (+3.97) 70.22 (+1.75)
4.1 Benchmarks
For mathematical reasoning experiments, we train models on 7.5k training samples in MATH (Hendrycks et al., 2021) and use MATH500 (Lightman et al., 2023) as the in-distribution test dataset. Additionally, we test the optimized models on out-of-distribution datasets: GSM8K (Cobbe et al., 2021), AIME24111https://huggingface.co/datasets/AI-MO/aimo-validation-aime, AMC23222https://huggingface.co/datasets/AI-MO/aimo-validation-amc, GaoKao2023En (Zhang et al., 2023), Minerva Math (Lewkowycz et al., 2022), and Olympiad Bench (He et al., 2024).
For LLM-as-a-Judge benchmarks, we train models on RewardBench (Lambert et al., 2024). Specifically, we convert the original data into a pair-ranking format with two possible orders, creating 5.97k items. We split them into a training set of 5k items and a test set of 970 items, denoted as RewardBench970. The models are also tested on JudgeBench (Tan et al., 2024) to assess out-of-distribution performance. We’ve tested two different split strategies, and we refer to Appendix B for detailed comparisons and results.
4.2 Baselines, Models, and Training Details
We compare performance across the following methods:
  • VRP (CoT): Prompting the LLM to solve the problem step-by-step (VRP in Sec 3.1).
[…]
  • ReMA (ours): Using RL to optimize the LLMs’ under MAMRP as defined in Eq 7.
And for each method, we show the greedy decoding performance in Section 4.3. We train and test Llama-3-8B-Instruct, Llama-3.1-8B-Instruct (Dubey et al., 2024), and Qwen2.5-7B-Instruct (Team, 2024) on mathematical reasoning benchmarks. For LLM-as-a-judge benchmarks, we train and test Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct. We use instruct-tuned LLMs so as to prompt them to perform VRP, MRP and MAMRP during training. Unless specified, we use two copies of the same model for high- and low-level agent in ReMA. We use the base reward setting in Sec E by default. For the underline RL algorithm, we use REINFORCE++ (Hu, 2025a). We refer to Appendix I for detailed training hyper-parameters and training settings.
4.3 Main Results
Table 1 compares the greedy decoding performance of ReMA against various RL baselines across mathematical reasoning benchmarks (Table 1(a)) and LLM-as-a-Judge benchmarks (Table 1(b)). Results across different LLMs indicate that, on average, ReMA outperforms all baselines, achieving a maximum average improvement of 6.68% on mathematical benchmarks and 8.49% on LLM-as-a-Judge benchmarks.
Notably, ReMA achieves the highest performance on most benchmarks, particularly on out-of-distribution datasets, with a maximum improvement of 20% on AMC23 for Llama3-8B-Instruct, 13.33% on AIME24 for Qwen2.5-7B-Instruct, 14.23% on RewardBench970 for Llama3.1-8B-Instruct. However, we observe that the accuracy gains from RL training on instruction-tuned models are smaller than from base models (Sec 4.4.1). Beyond the higher initial performance of instruction-tuned models, we hypothesize that this is due to their relatively fixed output distribution, which limits their adaptability and exploration in reinforcement learning. Additionally, the relatively small model size and dataset scale may further constrain the improvements achieved through reinforcement learning.
4.4 Ablations
4.4.1 Meta-thoughts boost low-level generalization
Figure 3: An RL experiment evaluates the performance of Qwen2.5-Math-7B on the MATH500, GSM8K, and AIME24 datasets after training on Level 3-5 MATH questions using different methods. RL from SFT achieves superior performance but struggles to generalize to more challenging problems. In contrast, RL from Base and RL under Meta-thinking demonstrate the ability to solve previously unseen, harder problems, with the latter further enhancing performance.
Here we provide a minimal but motivating example of how ReMA gives better learning dynamics. We use Qwen2.5-Math-7B (Yang et al., 2024) as the starting base model, MATH (level 3-5, about 5.5K number of instances) as the training dataset, and we compare three reinforcement learning training schemes, in particular:
  1. 1.
    RL from Base: train the base model directly on MATH with binary outcome reward;
  2. 2.
    RL from SFT: First, collect a small amount of CoT data (from GPT-4o’s answers) to fine-tune the base model (SFT); then train the tuned model on MATH with binary outcome reward;
  3. 3.
    RL under Meta-thinking: First, collect a small amount of high-level meta-thinking data (from GPT-4o’s plans) to fine-tune Qwen2.5-Math-7B-Instruct (Yang et al., 2024) to increase its meta-thinking abilities and make sure it act as high-level planner only; then train Qwen2.5-Math-7B as the low-level agent in collaboration with the high-level instruction with binary outcome reward.
The model is tested on a testing set of MATH500, GSM8K, and AIME24. The results are shown in Figure 3. First, supervised fine-tuning (SFT) largely boosts performance as it has the best accuracy on the in-distribution test set MATH500 and the easier one GSM8K at the beginning of RL training. However, it doesn’t show advances on the harder benchmark AIME24.
More importantly, while RL from SFT significantly boosts performance, it struggles to generalize to harder problems. In contrast, RL from Base shows limited improvements, especially on more challenging tasks. Notably, RL under Meta-thinking achieves the best learning dynamics, effectively leveraging high-level guidance to improve the performance of the lower-level solving agent, particularly on harder problems like AIME2024. We refer to Appendix J.1 for a detailed case study.
4.4.2 Reward functions shape cross-agent behaviors
We also investigate the impact of different reward function designs on ReMA’s behavior. In addition to the base reward setting described in Sec 3.3, we evaluate a consistency-based reward function using Qwen2.5-7B-Instruct. This reward function is designed to encourage the high-level agent to generate more detailed guidance. Indeed, we observe that the high-level agent trained in this manner produces more detailed solution steps compared to the one trained with the basic correctness format reward. However, we also find that this approach often leads to jailbreak behavior, where the high-level agent tends to include the final answer within its output, compromising the intended hierarchical reasoning process.
Furthermore, we discover an interesting evolution of a pattern during training: although our experimental setup is designed for the high-level agent to provide a solution plan while the lower-level agent executes it, we find that under the consistency-based reward, the lower-level agent significantly increases its attempt of verification rather than straightforward execution. We observed a certain sentence commonly appearing in the low-level agent’s responses: “Let’s go through the solution step by step to ensure clarity and correctness.” To quantify this effect, we track the frequency of it. We analyze this pattern across all mathematical test sets, sampling eight completions per question at a temperature of 0.7. Our empirical results have identified a 30x increase of such self-verifying patterns in the model trained with the consistency-based reward compared to the one trained with the base reward. Moreover, we also observe additional variations of this pattern, e.g. “Let’s carefully re-evaluate the problem and solution to ensure accuracy and clarity.” These phrases indicate that the low-level agent is actively exploring to verify the detailed response provided by the high-level agent.
This suggests that (1) meta-thinking can not only emerge and be reinforced in the high-level agent but also in the low-level agent. During reinforcement learning (RL) training, the two agents develop a novel problem-solving pattern characterized by a role reversal. (2) Consistency-based rewards promote a more self-corrective approach at the lower level, potentially disrupting the intended separation of roles between planning and execution. For a detailed case study, refer to Appendix J.2.
4.5 Diverse meta-thinking characteristics of LLMs
Figure 4: Average problem difficulty of three predefined metacognitive actions during training. Left: After approximately 20 training steps, the 1B LM’s outputs collapse to the simplest action, EMPTY, resulting in no data points for “REWRITE” and “DECOMPOSE” in the plot. Right: In contrast, the 8B LM learns to utilize more complex metacognitive actions for solving difficult problems.
To further analyze the meta-thinking behaviors learned by ReMA, we conduct a simple interpretability experiment through training with predefined JSON-format action sets inspired by Yue et al. . Specifically, we modify the high-level agent’s decision-making process by first selecting from three predefined meta-thinking actions: DECOMPOSE (breaking the problem into subproblems), REWRITE (reformulating the problem to simplify it), and EMPTY (solving the problem directly). After selecting an action, the agent then generates the corresponding text for that action within a single LM call, formatting the output in JSON.
We compare two experimental setups using Llama-3.1-8B-Instruct and Llama-3.2-1B-Instruct. These configurations allow us to study models of different scales: one setup consists of two 1B models, while the other consists of two 8B models. This setup enables us to investigate the impact of RL on the final training results of the high-level agent. To ensure the model maintains valid output formatting, we employ VLLM guided JSON decoding (Kwon et al., 2023; Dong et al., 2024) during rollout. We use the base reward setting, i.e., the averaged accuracy of the low-level agent’s solutions with format constraints.
First, we observe that the outputs of smaller language models (LMs) tend to converge to relatively simpler and shorter sentences compared to those in our main experiments. This may be due to the limited capacity of smaller LMs to generate valid JSON-formatted responses while simultaneously exploring diverse reasoning strategies. As shown in Fig. 4, Llama-3.2-1B-Instruct struggles to capture the complex dynamics underlying the task. Instead, it quickly converges to outputting only the EMPTY action to minimize penalties associated with formatting errors. In contrast, Llama-3.1-8B-Instruct successfully learns to employ different metacognitive strategies based on the difficulty level of the problem. A detailed case study is provided in Appendix J.3.
5 Related work
Drawing from the bitter lesson (Sutton, 2019), two methods that appear to scale effectively are searching and learning, aligning with current trends in large language models (Xu et al., 2025). At present, researchers are leveraging these methods to maximize the capabilities of individual transformers, while other efforts are exploring architectures that involve multiple interacting entities. In this paper, we examine this divergence within the context of LLM reasoning, a capability that allows large language models to solve problems through logical reasoning, step-by-step analysis, and inference (Wang et al., 2024a).
5.1 Single LLM Reasoning
Main research works in reasoning involving a single LLM utilize search-based and post-training methods. The fundamental elements of searching methods are text generation and evaluation. Generation schemes include In-Context Learning (Brown et al., 2020), Beam Search (Graves, 2012), and various tree-based searching (Snell et al., 2024); Evaluation approaches often use outcome accuracy, self-consistency (Wang et al., 2022), or process reward signal (Lightman et al., 2023) as the criteria to select high-quality responses from the generated texts. Post-training method is another research line in opposition to pre-training. Popular training pipelines often involve specific data construction followed by Supervised Fine-tuning (Qin et al., 2024; Ouyang et al., 2022; Hui et al., 2024; Liu et al., 2024), or reinforcement learning to interactively explore learning patterns (Wang et al., 2024a; Zhang et al., 2024a; DeepSeek-AI et al., 2025; Xu et al., 2025).
5.2 Multiple LLM Reasoning
Integrating multiple entities can potentially surpass the intelligence of the individual model (Chen et al., 2023). With the rapid emergence of large language models showing a varying level of abilities, some studies have explored facilitating discussions among multiple off-the-shelf LLMs (Zhang et al., 2025a; Chen et al., 2024a; Wang et al., 2023; Du et al., 2023; Zhuge et al., 2023; Tang et al., 2023; Hao et al., 2025; Akata et al., 2023; Hong et al., 2023; Zhang et al., 2024b), taking the form of free discussion (Du et al., 2023; Liang et al., 2023) or structured role assignments (Hong et al., 2023; Zhang et al., 2024b). Some have applied routing mechanisms to assign tasks to the most suitable expert models (Hu et al., 2024b; Stripelis et al., 2024; Ding et al., 2024; Yue et al., 2025; Chen et al., 2024c) or merging mechanisms to develop more versatile models (Yadav et al., 2024; Yu et al., 2024; Zhang et al., 2025b). Beyond aggregating static knowledge from multiple agents, multi-agent LLM training can also enhance reasoning capabilities. For example, multi-agent debates can generate diverse synthetic data, which can subsequently be used for supervised fine-tuning (Estornell et al., 2024; Li et al., 2024; Motwani et al., 2024; Dong & Ma, 2025; Perez et al., 2022; Ye et al., 2025a; Subramaniam et al., 2025). Reinforcement learning (RL) methods have also been adopted to improve LLM reasoning in areas such as alignment (Perez et al., 2022; Ma et al., 2023) and legibility (Kirchner et al., 2024a). Motwani et al. (2024) utilize a three-agent system for generation and fine-tune the models using Direct Preference Optimization (DPO). Reinforcement Learning with Generative Reward Models (GenRM) (Mahan et al., 2024; Ye et al., 2025b; Jiao et al., 2024; Wang et al., 2024b) represents another common approach of multi-agent training, where the reward signal is derived from the token probabilities of another LLM, coupled with the reasoning process. While our work aligns with these efforts, it diverges by using an additional tunable LLM to provide metacognitive instructions, guiding the low-level LLM during learning, rather than relying on a static GenRM. The most closely related works to ours are MAPoRL (Park et al., 2025) and COPRY (Ma et al., 2024). MAPoRL is a multi-agent debating framework that uses multi-agent reinforcement learning (MARL) with a learned verifier to fine-tune each LLM agent. COPRY duplicates an LLM into two agents, training them simultaneously in the roles of pioneer and observer using RL. However, unlike our approach, which explicitly separates metacognition from plan execution, these methods do not decompose the reasoning process but instead focus on improving direct chain-of-thought generation. Furthermore, our experiments are conducted on a larger scale and include more challenging problems.
5.3 Hierarchical Reasoning
Partitioning reasoning into hierarchical processes has been explored in prior research to make biological sense (Ye et al., 2018; Langley et al., 2004). In the context of language models, a hierarchical structure has been used to facilitate diverse reasoning patterns, including planning (Puerta-Merino et al., 2025; Sun et al., 2024; Song et al., 2023; Rana et al., 2023; Chen et al., 2024d; Yan et al., 2023; Xiao et al., 2024), validation (Haji et al., 2024; Xi et al., 2024) and self-refinement (Madaan et al., 2023; Kumar et al., 2024; Welleck et al., 2022). For instance, EvalPlanner (Saha et al., 2025b) is a framework that conducts reasoning through plan generation and execution. DOTS (Yue et al., 2024) extends decomposition by integrating a tree-based searching method with Analysis, Solution, and Verification layers. Marco-o1 (Zhao et al., 2024) focuses on open-ended problem-solving and abstract thinking, dynamically adjusting reasoning granularity and incorporating reflection mechanisms to enhance reasoning performance. Beyond these approaches, metacognition (Flavell, 1979) has been identified as another critical component of reasoning, referring to the intuitive understanding of one’s own cognitive and reasoning processes (Gao et al., 2024; Wang & Zhao, 2023). Wang & Zhao (2023) proposed a metacognitive prompting strategy to improve large language model (LLM) capabilities. Didolkar et al. (2024) further developed a prompt-guided method that enables models to label math problems with the required skills and subsequently use these labels to solve new problems. Gao et al. (2024) introduce meta-reasoner which use contextual multi-arm bandit to learn a high-level “advisor” over low-level reasoning process. Xiang et al. (2025) provides a Meta-CoT framework to think about its own thinking. They use construction-based methods as well as reinforcement learning to develop meta-cognitive skills. In our work, we also value reflecting on reasoning processes, and we enhance metacognitive abilities through two-agent interaction and reinforcement learning at both end.
6 Conclusion
In this paper, we propose ReMA, a novel framework that elicits meta-thinking in large language models (LLMs) using multi-agent reinforcement learning. Our approach explicitly separates meta-thinking from reasoning process by assigning them to distinct high- and low-level agents. This separation enhances exploration during training and improves result interpretability. We also introduce an iterative RL training procedure tailored to our framework. We evaluate ReMA on mathematical reasoning and LLM-as-a-Judge benchmarks, testing across different base models. To ensure robustness, we design two reward mechanisms: correctness-based and consistency-based rewards. Our results demonstrate that ReMA achieves the highest performance, particularly on out-of-distribution datasets. Additionally, we provide a case study illustrating how our collaborative training strategy enables effective coordination between high- and low-level models. The success of ReMA highlights the potential of reinforcement learning and the benefits of separating metacognition from reasoning execution.
7 Limitation and Future work
This work has several limitations that open opportunities for future research:
  • Our experiments were conducted exclusively on small language models (fewer than 8B parameters), which may not fully capture the potential benefits of scaling to larger models and datasets. We plan to extend our experiments to larger language models and more extensive datasets to better evaluate scalability and performance gains.
  • The proposed MRP/MAMRP methods did not consistently outperform the VRP baseline on the tested large language models. Additionally, our formulations are currently limited to single-turn interactions. We aim to refine these methods and extend them to multi-turn settings while investigating performance variations across different model architectures.
  • Our model has not yet been evaluated on a broad range of tasks, and we have not assessed the reasoning tax, i.e. the overthinking problem in current reasoning models. A comprehensive evaluation across diverse benchmarks and tasks is necessary, including a detailed analysis of the reasoning tax to understand its impact on overall performance. And the instruction of an extra meta-thinking agent shows great potential as a regulation to prevent over-thinking.
  • The current evaluation is restricted to mathematical tasks and the LLM-as-a-judge benchmark, which may not fully represent real-world scenarios. Future studies will include systematic comparisons across multiple benchmarks to better understand relative performance and identify specific areas for improvement.
We are actively working to improve the scalability of our MARL system during both training-time RL and test-time inference, ensuring that future iterations are more robust and broadly applicable.
Appendix A Inference-time Scaling of multi-agent metacognition-integrated reasoning process
In this section, we discuss how to enhance the inference-time computation of our hierarchical system, specifically focusing on the interaction between the high-level and low-level agents. The total number of model samples required for inference is determined by the product of the sampling budget allocated to each agent.
For instance, in a simple single-turn setting, if the high-level agent samples $k_{1}$ responses and each of these responses leads to $k_{2}$ samples from the low-level agent, the total number of model calls required is:
Total $\text{Total samples}=k_{1}\times k_{2}.$
Given a fixed computational budget, an important question arises: how should the sampling budget be distributed between the high-level and low-level agents to maximize performance? Allocating more samples to the high-level agent may increase diversity in reasoning strategies while allocating more to the low-level agent may yield more refined solutions for a given metacognitive plan.
Another crucial consideration is how to perform reranking on the final outputs. Two potential strategies include:
  • Hierarchical reranking: First, for each high-level response, rank and aggregate the low-level responses under it. Then, rank the aggregated results across different high-level responses.
  • Flat reranking: Directly rank all sampled responses together, regardless of the hierarchy of high-level reasoning steps.
Balancing sampling allocation and designing an effective reranking strategy are key challenges in efficiently scaling our multi-agent reasoning system. In the next section, we explore empirical results comparing different allocation strategies and ranking methods.
Appendix B Dataset Curation of RewardBench970
Table 2: Performance on LLM-as-a-Judge benchmarks, trained on dataset under the loose setting. The two-agent workflow in ReMA
[…]
Llama3.1 -8B -Instruct RewardBench970 71.24 81.86 (+10.62) 80.41 (+9.17) 86.29 (+15.05)
JudgeBench 51.77 51.45 (-0.32) 50.65 (-1.12) 53.71 (+1.94)
Average 61.51 66.65 (+5.14) 65.53 (+4.02) 70.00 (+8.49)
Qwen2.5 -7B -Instruct RewardBench970 86.49 87.22 (+0.73) 80.31 (-6.18) 90.72 (+4.23)
JudgeBench 58.39 54.84 (-3.55) 55.81 (-2.58) 58.71 (+0.32)
Average 72.44 71.03 (-1.41) 68.06 (-4.38) 74.72 (+2.28)
We process the original dataset in RewardBench by splitting it into a training set containing 5,000 tuples of (instruction, response A, response B) and a test set with the remaining 970 tuples.
To ensure a meaningful dataset split, we validate two separation strategies:
  • Loose setting: We only ensure that there is no direct overlap of tuples between the training and test sets.
  • Strict setting: We further enforce that no instruction appears in both the training and test sets. The results for this setting are presented in the main results (Table 1(b)).
Additionally, since the original RewardBench data originates from different subsets, we ensure that all original subsets are evenly represented in both the training and test sets.
Table 2 reports the learning performance of various methods under the loose dataset split setting. Compared to the results in Table 1(b), ReMA significantly outperforms other RL tuning baselines across all models, particularly on out-of-distribution (OOD) benchmarks. The consistent improvements on OOD datasets of these two settings suggest that ReMA enhances meta-thinking ability, resulting in better generalization across diverse task distributions.
Appendix C Improvement by RL Training
In this section, we illustrate the performance differences before and after RL training. We compare three inference settings:
  • VRP: A single LLM directly solves the problem step-by-step.
  • MRP: A single LLM first generates a meta-thinking message, and then solves the problem based on it.
  • MAMRP: One LLM generates the meta-thinking message, while another LLM follows it to solve the problem.
Figure 5: Relative performance gain after RL training for three different inference settings.
Figure 5 presents the results. We first evaluate these settings without RL training, i.e., using direct inference as a baseline. Then, we compare their performance with RL-trained models from Section 4.3. The results show that RL training enhances overall performance across all settings, with particularly notable improvements in MRP and MAMRP. This underscores the importance of RL fine-tuning in optimizing inference quality.
Appendix D Brief convergence analysis
Figure 6: Our method can be viewed as a combination of practical TRPO and block coordinate ascent, with the high and low-level models treated as distinct components within a larger neural network. Note that the figure does not represent the exact gradient back-propagation flow but rather highlights the key idea that we separate the high- and low-level models. This separation allows for the independent computation of gradients and the independent training of each model.
We reuse the notations from Section 3.2, where $\mathbf{x}$ is task prompt, $\mathbf{y}$ is generated answer, $\mathbf{y}^{*}$ is ground-truth, $\mathbf{m}$ is metacognition on task solving, $\pi_{\theta_{h}}$ and $\pi_{\theta_{l}}$ are high- and low-level agents with parameters $\theta_{h}$ and $\theta_{l}$. We consider the joint hierarchical policy defined in Equation 8 and update the objective as in Equation 9.
To leverage existing RL and optimization convergence analysis methods, we treat the two models as components of a larger model, as illustrated in Fig. 6. When updating one model, we treat the other model as part of a stationary environment. The gradients with respect to $\theta_{h}$ and $\theta_{l}$ are:
$\nabla_{\theta_{h}}J(\theta_{h},\theta_{l})=\mathbb{E}_{\mathbf{x},\mathbf{y}^{*}}\sum_{\mathbf{m}\sim\pi_{h}(\mathbf{m}\mid\mathbf{x};\theta_{h})}\nabla_{\theta_{h}}\pi_{h}(\mathbf{m}\mid\mathbf{x};\theta_{h})\left[\mathbb{E}_{\mathbf{y}\sim\pi_{l}(\mathbf{y}\mid\mathbf{x},\mathbf{m})}R(\mathbf{y},\mathbf{y}*)\right],$
$\nabla_{\theta_{l}}J(\theta_{h},\theta_{l})=\mathbb{E}_{\mathbf{x},\mathbf{y}^{*}}\sum_{\mathbf{y}\sim\pi(\theta_{h},\theta_{l})}\nabla_{\theta_{l}}\pi_{l}(\mathbf{y}\mid\mathbf{x},\mathbf{m};\theta_{h});\theta_{l})R(\mathbf{y},\mathbf{y}*).$
We can compute the gradients with log trick and estimate $\mathbb{E}_{\mathbf{y}\sim\pi_{l}(\mathbf{y}\mid\mathbf{x},\mathbf{m})}R(\mathbf{y},\mathbf{y}*)$ with Monte Carlo method.
Equipped with the objective function and gradient computation, we update the models iteratively. Without loss of generality, we analyze the case where the high-level policy is updated first:
$\theta_{h}^{(t+1)}=\arg\max_{\theta_{h}}J(\theta_{h},\theta_{l}^{(t)}),$
$\theta_{l}^{(t+1)}=\arg\max_{\theta_{l}}J(\theta_{h}^{(t+1)},\theta_{l}).$
Regarding the different regularizations $R_{h}$ and $R_{l}$ in Eq. 10 and 11 for the different policies, instead of directly integrating them into the loss function, we treat them as constraints, as done in Trust Region Policy Optimization (TRPO) (Schulman et al., 2015). Note that when one policy is fixed, the other policy operates in a stationary decision process.
Based on the defined objective and update method, we apply TRPO and block coordinate ascent to demonstrate that our update method theoretically converges to the optimal policies with respect to the final reward. First, recall that when updating a single policy, TRPO guarantees monotonic improvement by optimizing a lower bound. Specifically, let $\pi_{\text{old}}$ and $\pi$ represent the old and current policies, respectively. We define a surrogate objective as:
$L_{\pi_{\text{old}}}(\pi)=\mathbb{E}_{s\sim\pi_{\text{old}},a\sim\pi_{\text{old}}}\left[\frac{\pi(a|s)}{\pi_{\text{old}}(a|s)}A^{\pi_{\text{old}}}(s,a)\right],$
As shown by Schulman et al. (2015), the true objective of $\pi$ is lower-bounded by:
$J(\pi)\geq L_{\pi_{\text{old}}}(\pi)-C\cdot\max_{s}\operatorname{KL}[\pi_{\text{old}}(\cdot\mid s),\pi(\cdot\mid s)],$
for some constant $C$. By optimizing the right-hand side of the above inequality, we are guaranteed to improve the performance of $\pi$. Therefore, for policies $\pi^{t}$ and $\pi^{t+1}$ obtained from iterations $t$ and $t+1$ using the TRPO method, we have:
$J(\pi^{t+1})\geq J(\pi^{t}).$
Now, returning to our updating method, we treat the high- and low-level policies as two blocks of a single agent. The iterative update process can thus be viewed as a cyclic block coordinate ascent, where the two policies are updated in a fixed order. By updating each block using the TRPO method, and improving the surrogate objective within the KL constraint, each block update does not decrease $J$:
$J(\theta_{h}^{t+1},\theta_{l}^{t})\geq J(\theta_{h}^{t},\theta_{l}^{t}),$
$J(\theta_{h}^{t+1},\theta_{l}^{t+1})\geq J(\theta_{h}^{t+1},\theta_{l}^{t}).$
Thus $J(\theta_{h}^{t+1},\theta_{l}^{t+1})\geq J(\theta_{h}^{t},\theta_{l}^{t})$. This repeated coordinate maximization converges to a stationary point, where no single coordinate update can further improve $J(\theta_{h},\theta_{l})$.
Given the theoretical monotonic improvement with TRPO and block coordinate ascent, we adopt a practical version of TRPO in our experiments, specifically Proximal Policy Optimization (PPO) (Schulman et al., 2017) or REINFORCE++ (Hu, 2025b).
Appendix E Detailed reward design
To examine multi-agent metacognition-integrated reasoning with different reward designs, we experiment with different reward function designs to encourage effective collaboration and structured reasoning. Below, we introduce and justify several reward schemes.
1. Correctness and Format-Aware Reward (Base Setting)
In our primary reward setting, the system’s overall correctness is used as the primary reward signal, supplemented by format-based rewards for both the high-level and low-level agents. Using mathematical problem-solving as an example:
  • Low-level agent ($\pi_{\theta_{l}}$): Receives a reward of $+1.0$ for a correct answer. If the answer is incorrect, the agent is further penalized based on format compliance. Specifically:
    • If the output contains the designated answer-indicating format (e.g., boxed in LaTeX), it receives $-0.5$.
    • Otherwise, it receives $-1.0$, as a missing format often suggests an incomplete or unstructured response.
  • High-level agent ($\pi_{\theta_{h}}$): Receives the average correctness of the low-level agent’s sampled responses as its reward. Additionally, to prevent the high-level agent from directly generating explicit answers instead of guiding reasoning, a strong penalty of $-1.0$ is applied if it includes an explicit answer format (e.g., boxed).
2. Consistency-Based Reward
Instead of using correctness as the high-level reward signal, this approach rewards the high-level agent for promoting consistent responses from the low-level agent, regardless of actual correctness. The consistency reward is defined as the proportion of the most frequently occurring answer among all sampled responses:
$R_{h}=\frac{\text{max occurrence of an answer}}{\text{total number of responses}}$ (12)
If the majority of responses do not contain a definitive answer, the reward is set to zero. We also add the format penalty to the high-level agent if its output contains the designated answer-indicating format. This incentivizes the high-level agent to guide the low-level agent toward more stable, detailed, reproducible outputs rather than erratic reasoning paths.
These different reward formulations allow us to investigate various dimensions of metacognitive reasoning: correctness, consistency, etc. We empirically compare their effects on learned metacognitive reasoning patterns in Sec 4.4.2.
Appendix F Learning to reason from the perspective of Leader Follower Game
Besides the loss function in the main part, we also propose to frame the problem as a leader-follower game. By analyzing the equilibria of the leader-follower game, we demonstrate that our framework inherently identifies the optimal sub-tasks aligned with the capabilities of the low-level model. This ensures that the high-level decisions are guided by the low-level model’s strengths, leading to more efficient and targeted task decomposition.
F.1 Leader-follower game
The leader-follower game, also known as the Stackelberg game, models interaction between two agents with parametrized strategies $\bm{w}=(\bm{w}_{1},\bm{w}_{2})$ and differentiable objective functions $(\mathcal{L}_{1},\mathcal{L}_{2}):\mathbb{R}^{d}\rightarrow\mathbb{R}$. In this framework, the leader announces its strategy first, and the follower observes this decision to respond optimally. This sequential structure enables the leader to anticipate the follower’s reaction and adjust its strategy accordingly. A Stackelberg equilibrium occurs when neither agent can unilaterally improve its objective. Denoting $\bm{w}_{1}$ as the leader’s strategy and $\bm{w}_{2}$ as the follower’s, the loss functions $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are optimized with the following bi-level structure:
$\bm{w}_{1}^{*}=\text{argmin}_{\bm{w}_{1}}\mathcal{L}_{1}(\bm{w},\bm{w}_{2}^{*}(\bm{w}_{1})),\quad\bm{w}_{2}^{*}(\bm{w}_{1})=\text{argmin}_{\bm{w}_{2}}\mathcal{L}_{2}(\bm{w}_{1},\bm{w}_{2}).$
Anil et al. (2021) apply the leader-follower game to ensure checkable answers in a prover-verifier game (PVG). The objective is a verifier that is both complete (accepts all correct proofs from a verifier) and sound (rejects all incorrect proofs from a verifier). They analyze different scenarios where the verifier acts as the leader, the prover as the follower, and both announce strategies simultaneously, forming a Nash equilibrium. The study concludes that in verifier-led PVG, a Stackelberg equilibrium is both necessary and sufficient for achieving a sound and complete verifier, whereas in other configurations, a Stackelberg equilibrium is not necessary or sufficient for this outcome.
F.2 Efficacy of LLM
Because the high-level policy possesses strong generalization capabilities, it is impractical for it to exhaustively explore every potential sub-task for each question. Instead, it naturally focuses on tasks within a feasible range of difficulty, leveraging only a limited set of coarse planning actions. Rather than pinpointing perfectly tailored sub-tasks, the policy searches for general tasks of particular computational complexity, i.e., difficulty, that it can handle reliably. Motivated by this perspective, we incorporate the concept of a reasoning boundary for large language models (LLMs) (Chen et al., 2024b). Intuitively, the reasoning boundary circumscribes the maximum difficulty of problems a model can solve at a desired accuracy level. Formally, for a model $\theta$, a task $t$, and a predefined threshold $A$, the reasoning boundary of $\theta$ represents the maximum problem difficulty $d$ that satisfies:
$\mathcal{B}_{Acc=A}(t|\theta)=\sup_{d}\{d|Acc(t|d,\theta)=A\}.$
where $d$ denotes the problem difficulty. By quantifying the difficulty level a model can reliably handle, the reasoning boundary provides a systematic way to align the high-level policy’s focus with the model’s actual capabilities, gauge the efficacy of the low-level policy, and determine the optimal strategy for solving the question.
F.3 Leader-follower Game for LLM Reasoning
Our goal is to find the high-level policy that searches for the sub-task sequence based on the efficacy of the low-level policy to solve the question. We design the loss functions as follows:
$\mathcal{L}_{h}=\mathbb{E}_{(x,y)\sim p_{D},t_{1:K}}\left[-\log\pi_{l}(y_{K}\mid x,t_{1:K},y_{1:K-1})\right],$
$\mathcal{L}_{l}=\mathbb{E}_{x\sim p_{D},t_{1:k}\sim\pi_{h},\hat{y}_{k}\sim\pi_{l}}\left[-r(y_{k},\hat{y}_{k}\mid x,t_{1:k},y_{1:k-1})\right],$
where $r(y_{k},\hat{y}_{k}\mid x,t_{1:k},y_{1:k-1})$ represents the step reward for the correctness of $\hat{y}_{k}$ derived from the question $x$, the sub-task sequence $t_{1:k}$ from the high policy and prior intermediate answer $y_{1:k-1}$. The loss functions can be interpreted as follows: the high-level policy is incentivized to find sub-tasks that lead to the correct answer based on the capabilities of the low-level policy, while the low-level policy is incentivized to enhance its instruction-following ability.
How to minimize the loss functions and whether such minimization leads to the desired results remain questions. To explore this, we consider a simplified case of our method, where the high-level policy plans the complete sub-task sequence at the beginning and the low-level executes the instruction in a single interaction. The corresponding parameterized policies are defined as $\pi_{h}((t_{1},\dots,t_{K})\mid x)$ and $\pi_{l}((\hat{y}_{1},\dots,\hat{y}_{K})\mid x,(t_{1},\dots,t_{K}))$. The corresponding loss functions are:
$\mathcal{L}_{h}=\mathbb{E}_{(x,y)\sim p_{D},t_{1:K}}\left[-\log\pi_{l}(y_{K}\mid x,t_{1:K})\right],$ (13)
$\mathcal{L}_{l}=\mathbb{E}_{x\sim p_{D},t_{1:k}\sim\pi_{h},\hat{y}_{k}\sim\pi_{l}}\left[-r(y_{k},\hat{y}_{k}\mid x,t_{1:k},y_{1:k-1})\right].$ (14)
In this step, the high-level policy generates the entire sub-task sequence without relying on intermediate answers, while the low-level policy follows the sequence to produce answers for the sub-tasks. The low-level policy can still leverage prior intermediate answers to sequentially refine its responses.
To analyze the result agents by minimizing the loss functions, we adopt the completeness and soundness properties from the PVG framework for LLM reasoning. Specifically, if the high-level policy generates a sub-task sequence that is executable within the low-level policy’s capabilities, the problem must be solved (completeness). Conversely, if the sub-task sequence is incorrect or beyond the low-level policy’s capacity, the problem cannot be solved (soundness). To achieve this, we utilize the conclusion from Anil et al. (2021), which positions the low-level policy as the leader and the high-level policy as the follower, equilibria guarantee the complete and sound low-level policy.
When the high-level policy takes the lead, the low-level policy is forced to adapt to the specific strategy defined by the high-level policy, which can result in neither complete nor sound low-level policy. For example, if the high-level policy dictates that it will only generate sub-tasks involving addition and subtraction, the low-level policy is constrained to optimize only for these tasks. While they may reach an equilibrium, the low-level policy remains incomplete, and this limitation impacts both policies. In the case of the simultaneous PVG game, convergence to a Nash equilibrium is possible, but it is not sufficient for completeness and soundness. For instance, the low-level policy might disregard the high-level policy entirely (e.g., if the high-level provides incorrect instructions, but the low-level still performs correctly). This approach, however, is challenging to implement due to the significantly larger search space involved.
Furthermore, the loss functions we design ensure that, at a Stackelberg equilibrium, the high-level policy identifies sub-task sequences that the low-level policy can execute to solve the problem with the highest probability. With the low-level policy acting as the leader, it establishes its reasoning boundary for tasks. Based on the reasoning boundary, let $\theta_{h}$ and $\theta_{l}$ represent the policy parameters for the high-level and low-level policies, respectively. The probability that the low-level policy correctly solves the question is defined as:
$\pi_{l}(y_{K}\mid x,t_{1:K})=\prod_{k=1}^{K}\text{Acc}(t_{k}\mid x,\theta_{l}),$
where we can compute the difficulty $d_{k}$ from $t_{k}$ and $x$. where the difficulty $d_{k}$ can be derived from $t_{k}$ and $x$. The loss function in Equation 13 ensures that the selected sub-tasks are optimal for the low-level policy. Here we provide a theoretical condition under which the most efficient solution strategy can be identified, according to the efficacy of the LLM.
This approach can be viewed as a game between a high-level ”prover” and a low-level ”verifier”. The verifier, representing the low-level policy, adheres the high-level policy’s instructions to validate its reasoning. Unlike the classic PVG setting, where the prover has ground-truth labels, the label of our high-level policy depends on the tunable low-level policy. This distinction, where the low-level policy (leader) is inherently more complex, contrasts with traditional PVG setups and adds complexity due to the interdependence between the high- and low-level policies.
By framing the problem-solving process as a leader-follower game, with the low-level policy designated as the leader, we can construct a bi-level optimization problem to identify an equilibrium. Following the formulation in Section F.1, the problem is expressed as:
$\theta_{l}^{*}=\operatorname*{arg\,min}_{\theta_{l}}\mathcal{L}_{l}(\theta_{h}^{*}(\theta_{l}),\theta_{l})\quad\theta_{h}^{*}(\theta_{l})=\operatorname*{arg\,min}_{\theta_{l}}\mathcal{L}_{h}(\theta_{h},\theta_{l}).$
Then we can apply bi-level optimization techniques.
Appendix G Pseudocode of ReMA
The pseudocode is shown in Algorithm 1.
Algorithm 1 Single turn MAMRP
0:  High-level policy $\pi_{h}$, Low-level policy $\pi_{l}$, Dataset $\mathcal{D}$, Optimizers for $\pi_{h}$ and $\pi_{l}$. $\varepsilon_{\min},\varepsilon_{\max}$ to filter training dataset
1:  Initialize $\pi_{h}$ and $\pi_{l}$
2:  while not converged do
3:     build training dataset $\mathcal{D}_{l}$ with $\pi_{h},\pi_{l},\varepsilon_{\min},\varepsilon_{\max}$
4:     for Sample $(\mathbf{x,m},\mathbf{y^{*}})\sim\mathcal{D}_{l}$ do
5:        Generate $\mathbf{y}\sim\pi_{l}(\mathbf{x},\mathbf{m})$
6:        Compute low-level reward $R_{l}(\mathbf{y},\mathbf{y^{*}})$
7:        Update $\pi_{l}$ using $\nabla_{\theta_{l}}\mathbb{E}[R_{l}]$
8:     end for
9:     build training dataset $\mathcal{D}_{h}$ with $\pi_{h},\pi_{l},\varepsilon_{\min},\varepsilon_{\max}$
10:     for Sample $(\mathbf{x},\mathbf{y^{*}})\sim\mathcal{D}_{h}$ do
11:        Generate $\mathbf{m}\sim\pi_{h}(\mathbf{x})$ and $\mathbf{y}\sim\pi_{l}(\mathbf{x},\mathbf{m})$
12:        Compute high-level reward $R_{h}(\mathbf{m},\mathbf{y},\mathbf{y^{*}})$
13:        Update $\pi_{h}$ using $\nabla_{\theta_{h}}\mathbb{E}[R_{h}]$
14:     end for
15:  end while
Appendix H Supervised fine-tuning data collection
For experiments in Sec 4.4.1, we collect expert data to enhance the reasoning pattern, i.e. RL from SFT. Specifically, we collect demonstration data from GPT-4o Mini on MATH training dataset (7.5k problems) Hendrycks et al. (2021) and use it to fine-tune the LLMs. The data generation follows these steps: First, we prompt GPT-4o Mini to produce metacognitive reasoning for high-level model training. Specifically, we use different prompts to instruct it to rewrite and decompose a given question without providing a final answer. We collect metacognitive reasoning using two predefined actions, “rewrite” and “decompose”, which align with human approaches to complex problem-solving while preserving answer diversity. Next, we use the generated instructions to prompt GPT-4o Mini to follow the metacognitive steps and solve the question, obtaining SFT data for low-level policy training. Below, we present the prompts used for both high-level and low-level models.
Prompt for metacognition reasoning to rewrite:
System prompt: You are a math expert trying to solve mathematical problems. Before answering a question, your task is to rewrite the original question to make it clearer. Provide your rewritten content in JSON format: {{ "action": "REWRITE", "output": "{{clearer question content}}" }} Respond only with valid JSON. Do not write an introduction or summary. User prompt: Here is the question: [problem_text]
Prompt for metacognition reasoning to decompose:
System prompt: You are a math expert trying to solve mathematical problems. Before answering a question, your task is to decompose the original question to make it clearer. Provide your rewritten content in JSON format: {{ "action": "DECOMPOSE", "output": "{{decomposed question content}}" }} Respond only with valid JSON. Do not write an introduction or summary. User prompt: Here is the question: [problem_text]
Prompt for generating final answers using on the question and metacognition reasoning:
[…]
Appendix I Training Details
We refer to Appendix K for prompts we use during training. We implement the training pipeline with OpenRLHF (Hu et al., 2024a) which is a highly efficient codebase and is easy to scale up. We select REINFORCE++ to save resources and for efficient training. All experiments are conducted in a node of 8 NVIDIA A100 GPUs. We use bf16, Zero2, Flash-Attention and gradient checkpointing to run our experiments.
During rollout, we set temperature=1.0, top_p=1.0, top_k=-1, and use vLLM for inference acceleration. We set the max generation length to be 2048 and, the rollout batch size to be 1000. The number of samples per prompt is 4. During training, we use Adam Optimizer with a learning rate of 5e-7. We set the mini-batch size to be 500, and the clip ratio to be 0.2. Other hyperparameters, such as KL coefficients and the number of training episodes, were carefully tuned based on validation set performance to ensure robust and reliable results. To align with the hyperparameter in OpenRLHF, we use #Training Episode as the number of reinforcement learning epoch on the entire dataset.
In ReMA, during prompt filtering of the high-level model, the high-level agent first samples 10 candidates for each question with t=1.0, and for each output the low-level agents sample 1 solution with t=0.0, then we select questions of success rate between $[\varepsilon_{\min},\varepsilon_{\max}]$. And for the low-level agent’s prompt filtering, the high-level agent first samples 1 candidate for each question with t=0.0, and for each output the low-level agents sample 10 solutions with t=1.0, then we select questions of success rate between $[\varepsilon_{\min},\varepsilon_{\max}]$ and use the high-level agent to sample 4 meta-thoughts with t=1.0 as the input.
I.1 Training on MATH
VRP
For Llama3-8B-Instruct, Llama3.1-8B-Instruct, and Qwen2.5-7B-Instruct, we all use a KL coefficient of 1e-2, and for #Training Episode, we use 12,6,6 for these 3 models respectively. For Llama3-8B-Instruct, we set the learning rate of 2e-7 for stable training.
MRP
For Llama3-8B-Instruct, Llama3.1-8B-Instruct, and Qwen2.5-7B-Instruct, we all use a KL coefficient of 1e-2, and for #Training Episode, we use 10,6,6 for these 3 models respectively.
MAMRP
We use $\varepsilon_{\min}=0.2,\varepsilon_{\max}=0.8$ for prompt filtering. We use the same #Training Episode=4 for all models, and for #Update Iteration, we use 3 for Llama3-8B-Instruct and Llama3.1-8B-Instruct, 10 for Qwen2.5-7B-Instruct. And we set the KL coefficient to be 1e-2 for all the 3 models.
I.2 Training on Reward Bench
VRP
For Llama3.1-8B-Instruct, and Qwen2.5-7B-Instruct, we all use a KL coefficient of 1e-2, and for #Training Episode, we use 4,6 for these 2 models respectively.
MRP
For Llama3.1-8B-Instruct, and Qwen2.5-7B-Instruct, we all use a KL coefficient of 1e-2, and for #Training Episode, we use 4,6 for these 2 models respectively.
MAMRP
We set #Update Iteration=1 for all models. We set the KL coefficient to be 1e-2 for Llama3.1-8B-Instruct and 1e-2 for Qwen2.5-7B-Instruct all models. For Llama3.1-8B-Instruct, we use $\varepsilon_{\min}=0.2,\varepsilon_{\max}=0.8$ for prompt filtering and we use #Training Episode of 2 during training. For Llama3.1-8B-Instruct, we use $\varepsilon_{\min}=0.1,\varepsilon_{\max}=0.9$ for prompt filtering and we use #Training Episode of 1 during training.
Appendix J Qualitative Results
J.1 High-level policy finds better plans
Here is an example of how a high-level policy alters the solving method of an LLM, increasing the likelihood of providing correct answers. As we can see from the following example, without the high-level policy, the LLM counts all integer coordinates, including those on the boundary, and then subtracts the boundary coordinates. In contrast, the high-level policy identifies a better approach, directly instructing the LLM to count only the coordinates strictly inside the boundary.
Figure 7: Case Study comparing with and without high-level metacognition results.
J.2 Case study for Experiments in Sec 3.3
Fig 8 and Fig 9 show an case study of experiments in Sec 3.3.
Although both agents are prompted with the same instructions as in our main results, the consistency reward of the high-level agent significantly alters the learning dynamics. As illustrated in Fig. 8, the high-level agent generates detailed solution attempts rather than a strategic plan. Consequently, the low-level agent evolves to verify the high-level agent’s solutions. This suggests that, during reinforcement learning (RL) training, the two agents develop a novel problem-solving pattern characterized by a role reversal, where one agent generates an answer while the other verifies it.
Figure 8: Case Study for consistency reward of high-level agent
Figure 9: Case Study for base correctness reward of high-level agent
J.3 Case study for Experiments in Sec 4.5
Fig 10 shows the detailed outputs of the two groups of LLMs trained in JSON action. The outputs finally converge to relative easier and shorter sentences, compared to our main experiments. This may due to insufficient capacity of a small LM to produce a valid JSON-format response while simultaneously exploring diverse reasoning strategies.
Figure 10: Case Study for interpretability experiments in Sec 4.5
Appendix K Prompts
K.1 Prompts for Math problems
VRP prompt:
[…]
MRP prompt:
[…]
MAMRP prompt: high-level agent:
System prompt: You are a math expert specialized in solving mathematical problems, you need to teach a weaker agent with minimal capability in math how to solve a problem step-by-step. Your task is to provide a high-level solution plan for the given problem, in order to guide a low-level math solving agent to solve the problem. You can not directly answer the question. You’ll be punished if you include any answer in your response. You need to first think deeply in mind and output your final instruction. User prompt: Here is the question: {Question} low-level agent:
[…]
[Question] {Question} [End of Question]
[Provided Instruction] {instruction} [End of Instruction]
K.2 Prompts for LLM-as-a-Judge problems
We adopt the prompts from Saha et al. (2025a). VRP prompt:
System prompt: Please act as an impartial judge and evaluate the quality of the responses provided by two AI assistants to the user question displayed below. You should choose the assistant that follows the user’s instructions and answers the user’s question better. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of their responses. Begin your evaluation by comparing the two responses and provide a short explanation. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Do not favor certain names of the assistants. Be as objective as possible. After providing your explanation, output your final verdict by strictly following this format: "[[A]]" if assistant A is better, "[[B]]" if assistant B is better. User prompt: [User Question] {instruction} [End of User Question] [The Start of Assistant A’s Answer] {response_A} [The End of Assistant A’s Answer] [The Start of Assistant B’s Answer] {response_B} [The End of Assistant B’s Answer]
MRP prompt:
System prompt: Please act as an impartial judge and evaluate the quality of the responses provided by two AI assistants to the user question displayed below. You should choose the assistant that follows the user’s instructions and answers the user’s question better. First of your task is to build an evaluation plan that can then be executed to assess the response quality. Whenever appropriate, you can choose to also include a step-by-step reference answer as part of the evaluation plan. Enclose your evaluation plan between the tags "[Start of Evaluation Plan]" and "[End of Evaluation Plan]". After that, please act as an impartial judge and evaluate the quality of the responses provided by two AI assistants to the user question displayed below. You should choose the assistant that follows the user’s instructions and answers the user’s question better. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of their responses. Begin your evaluation by comparing the two responses and provide a short explanation. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Do not favor certain names of the assistants. Be as objective as possible. After providing your explanation, output your final verdict by strictly following this format: "[[A]]" if assistant A is better, "[[B]]" if assistant B is better. User prompt: [User Question] {instruction} [End of User Question] [The Start of Assistant A’s Answer] {response_A} [The End of Assistant A’s Answer] [The Start of Assistant B’s Answer] {response_B} [The End of Assistant B’s Answer]
MAMRP prompt: high-level agent:
System prompt: We want to evaluate the quality of the responses provided by AI assistants to the user question displayed below. For that, your task is to help us build an evaluation plan that can then be executed to assess the response quality. Whenever appropriate, you can choose to also include a step-by-step reference answer as part of the evaluation plan. User prompt: [User Question] {Question} [End of User Question]
low-level agent:
System prompt: Please act as an impartial judge and evaluate the quality of the responses provided by two AI assistants to the user question displayed below. Your evaluation should be performed by following the provided evaluation plan step-by-step. Avoid copying the plan when doing the evaluation. Please also only stick to the given plan and provide explanation of how the plan is executed to compare the two responses. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Do not favor certain names of the assistants. Be as objective as possible. After providing your evaluation, output your final verdict by strictly following this format: "[[A]]" if assistant A is better, "[[B]]" if assistant B is better. User prompt: [User Question] {instruction} [End of User Question] [The Start of Assistant A’s Answer] {response_A} [The End of Assistant A’s Answer] [The Start of Assistant B’s Answer] {response_B} [The End of Assistant B’s Answer] [The Start of Evaluation Plan] {evaluation_plan} [The End of Evaluation Plan]